OpenStudy (anonymous):

Please help! :) -5x+2y=4 -10x+y=-13

hero (hero):

Subtract the second equation from the first: -5x -(-10)x + 2y - y = 4 -(-13) -5x + 10x + y = 4 + 13 -5x + y = 17 y = 5x + 17 Now substitute that back into the second equation: -10x + (5x + 17) = -13 And solve for x.

OpenStudy (anonymous):

alright thanks!

hero (hero):

and solve for y after you find x.

OpenStudy (anonymous):

yea i got it thanks

hero (hero):

Did you solve for x and y? What did you get?

OpenStudy (anonymous):

x=2 and y=7